OpenStudy (anonymous):

Add and simplify:

OpenStudy (anonymous):

\[4 \over m+2 \] + \[2 \over m-2\]

OpenStudy (anonymous):

Answer choices: A. \[6 \over (m+2) (m-2)\] B. \[6m-4 \over (m+2) (m-2)\] C. \[6m+12 \over (m+2) (m-2)\]

OpenStudy (anonymous):

B

OpenStudy (anonymous):

B

OpenStudy (lalaly):

\[(\frac{4}{m+2} \times \frac{m-2}{m-2} )+( \frac{2}{m-2} \times \frac{m+2}{m+2})\]\[=\frac{4(m-2)+2(m+2)}{(m-2)(m+2)}= \frac{6m-4}{(m+2)(m-2)}\]

OpenStudy (anonymous):

B

OpenStudy (anonymous):